OpenStudy (anonymous):

Rewrite f(x) = –2(x − 3)^2 + 2 from vertex form to standard form.

OpenStudy (anonymous):

@GabbyLovesYou @Loser66 @Nnesha @welshfella

OpenStudy (welshfella):

expand the (x - 3)^2 first

OpenStudy (welshfella):

(x - 3)(x - 3) = ?

OpenStudy (anonymous):

x^2 - x^3 - x^3 + 9

OpenStudy (welshfella):

well you have the first and last terms correct but 3 times x is not x-cubed

OpenStudy (welshfella):

x * x * x = x^3 3 * x = 3x

OpenStudy (anonymous):

Ah, x^2 - 3x - 3x + 9

OpenStudy (welshfella):

right now you can add the 2 middle terms to give x^2 - 6x + 9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Multiply by -2?

OpenStudy (welshfella):

so we have -2(x^2 - 6x + 9) + 2 now distribute the -2 over the parentheses - which basically means you multiply (x^2 - 6x + 9) by -2

OpenStudy (anonymous):

-2x^2 +12x - 18

OpenStudy (welshfella):

right - but dont forget the + 2

OpenStudy (anonymous):

Ah, -2x^2 +12x - 20

OpenStudy (welshfella):

not quite - 18 + 2 = -16

OpenStudy (anonymous):

Oh, right, didn't see the "-".

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (welshfella):

yw